He decided that he wants to get the largest garden from 200 feet of fencing. What are the dimensions of his garden? What if he bought n feet of wire?
Found 2 solutions by checkley71, final:
Answer by checkley71(8403)

THE LARGEST AREA OF ANY RECTANGLE IS A SQUARE THUS:
4X=200
X=200/4
X=50 FEET FOR EACH SIDE.
THUS THE AREA=50*50=2,500 FT^2
IF HE BOUGHT N FEET OF WIRE THEN HIS GARDEN WILL BE:
[(N/4)^2] FT^2

Answer by final(2)

hey
i dont understand how you solved it.
can you explain how you got the equations